Buoyance

A rock measures on a scale at 235.0 grams ± 0.5 grams, when the rock is submersed in water the scale reads 200.0 ± 0.5 grams. Calculate the specific gravity of the rock, with uncertainty Calculate the volume of the rock, with uncertainty.

You dunk that rock into a mysterious green goo, the scale reads 220.0 grams, what is the density of the fluid? (no uncertainty necessary)
Answer #1

mass of the rock in air mo = 235 +/- 0.5 g

mass of water in water mw = 200 +/- 0.5 g

loss of weight in water is equal to weight of water displaced

Volume of water displaced = (mo - mw ) cc ( density of water = 1 gm/cc)

= volume of the rock

density of rock dr = mo / (mo - mw ) gm /cc

specific gravity of rock = dr /dw =  mo / (mo - mw )

= 235/(235-200) = 6.71

235 +/- 0.5 = 235 +/- 0.2 %

uncertainty in denominator = 0.5 +0.5 = 1

i.e 35 +/-1 = 35 +/- 2.86 %

uncertainty for SG = 0.2 % + 2.86% = 3.06 %

SG = 6.71 +/- 3.06 % = 6.71 +/- 0.21

volume of the rock = (235 +/-0.5) - (200 +/- 0.5)

= 35.0 +/- 1.0 gm/cc

when rock is sunk in goo loss of weight = Vr * dg ( dg density of goo)

= 235 - 220 = 15 g

density of goo = 15/35 = 0.43 gm/cc